# Partial Law Invariance and Risk Measures

### v4

@@
Abstract
We introduce the concept of partial law invariance, generalizing the concepts of law invariance and probabilistic sophistication widely used in decision theory, as well as statistical and financial applications. This new concept is motivated by practical considerations of decision making under uncertainty, thus connecting the literature on decision theory and that on financial risk management. We fully characterize partially law-invariant coherent risk measures via a novel representation formula. Strong partial law invariance is defined to bridge the gap between the above characterization and the classic representation formula of Kusuoka. We propose a few classes of new risk measures, including partially law-invariant versions of the Expected Shortfall and the entropic risk measures, and illustrate their applications in risk assessment under different types of uncertainty. We provide a tractable optimization formula for computing a class of partially law-invariant coherent risk measures and give a numerical example.
Keywords: Law invariance, probabilistic sophistication, ambiguity, coherent risk measures, source theory, Expected Shortfall
@@ 1 Introduction
Law invariance is the fundamental property that makes risk measures, as well as many decision models, applicable in practice. Once a reference probability model is specified, either through statistical estimation or expert opinion, law-invariant risk measures evaluate financial positions that share the same distribution under this probability model equally. This connects the theory of risk measures to statistics, thus allowing for a description of the probabilistic structure of financial positions, arguably a basic step in any quantitative financial analysis. In decision theory, law invariance is often known as probabilistic sophistication (Machina and Schmeidler (1992)). Although law invariance is a fundamental concept, it may not be universally desirable in all contexts. For instance, model misspecification, or model uncertainty, occurs when the assumed probabilistic model for various financial positions does not accurately represent the true, unknown, distribution of these positions. This can be problematic as the value of a law-invariant risk measure may differ significantly between the assumed model and the objective truth. For instance, during the 2008 financial crisis, the probability model used for default rates of mortgages was questionable, and the loss distribution for mortgage-backed securities was underestimated, leading to insufficient capital requirements for many companies to remain solvent (e.g., McNeil et al. (2015)). Although we take a risk measure perspective in this paper, the problem of model uncertainty extends beyond this and can affect any (usually statistical) process that relies on estimating a probabilistic model, followed by computing a quantity based on that model. We will generally refer to uncertainty in the probabilistic model as “model uncertainty”, which is also known as Knightian uncertainty or ambiguity in the literature. There are many avenues of research that account for model uncertainty in risk assessment in different ways. One area is distributionally robust optimization, in which optimization with risk measures is performed against a worst-case scenario; see e.g., Ghaoui et al. (2003), Natarajan et al. (2008), and Zhu and Fukushima (2009). Another area is to study risk measures formulated from multiple scenarios; see e.g., Delage and Li (2018), Delage et al. (2019) and Wang and Ziegel (2021). Model uncertainty is deeply rooted in decision theory (often called ambiguity there), with influential work pioneered by Schmeidler (1989) and Gilboa and Schmeidler (1989), followed by e.g., Klibanoff et al. (2005) and Maccheroni et al. (2006).
Most of the literature, including the works cited above, addresses model uncertainty through two approaches. The first approach begins with a reference probability model and a law-invariant risk measure. This risk measure would reflect an individual’s attitudes towards the financial positions if the individual had complete confidence in the reference probability model. Subsequently, to account for model uncertainty, the risk measure is aggregated over multiple alternative probability models, where each alternative model is motivated by the reference probability model. The aggregation is often, but not always, the worst-case aggregation, which is the approach in distributionally robust optimization, where the law-invariant risk measure is a commonly used one, such as VaR or ES. The second approach begins with a general risk measure that depends solely on the outcomes of the financial positions across possible future states of the world. Unlike the first approach, this approach does not assume a reference probability model, meaning the risk measure cannot be law invariant from the outset. This method circumvents model uncertainty by never supposing a model from the start, and it is consistent with the decision models of Schmeidler (1989) and Maccheroni et al. (2006). In the context of risk measures, this approach is quite common; see Föllmer and Schied (2016).
In this paper, we take a novel approach. We begin with a reference probability model but not a law-invariant risk measure. To account for model uncertainty, risk measures are assumed to satisfy law invariance only partially, in a sense to be specified later. The following example can motivate this idea. After estimating a probability model for various loss factors, it is natural that different factors will have different levels of model uncertainty. For example, this difference might result from a statistical inference procedure, where the confidence regions for parameters tied to certain loss factors could be larger or smaller than those for other loss factors, for instance, due to data scarcity. From this point, we can typically divide the loss factors into two categories: the loss factors for which we have no, or negligible, model uncertainty and those that have non-negligible model uncertainty. We will call all the events related to the loss factors in the first category the “model-certain source.” A financial position is said to depend on this source if its outcome is determined by events within this source. The next step is to select a risk measure for decision making. Requiring law invariance would not be able to distinguish factors in or outside the model-certain source, which is not desirable. However, for financial positions depending on the model-certain source, it is reasonable to measure their risk level solely based on distributions given by the probability model, as these positions involve no model uncertainty; thus, a partial form of law invariance is desirable. With this guiding principle in mind, this paper introduces a new property called partial law invariance. Fix a source, mathematically modeled by a sub-$\sigma$-algebra $\mathcal{G}$ of the whole model space. A risk measure is partially law invariant (with respect to $\mathcal{G}$) if it equates the risks of financial positions that (a) both belong to this source (i.e., $\mathcal{G}$-measurable) and (b) share identical distributions. This property reflects the discussion above with $\mathcal{G}$ being the model-certain source, but it is applicable beyond this interpretation. Although we take risk measures as the main objects of the paper, the concept of partial law invariance is naturally formulated for preference relations, applicable in broader contexts than financial risk assessment.
@@
The approach of partial law invariance to modeling risk measures can address situations where some sources of randomness are more uncertain than others. Traditional approaches, such as distributionally robust optimization, cannot handle this setting, as these models of uncertainty cannot directly incorporate sub-$\sigma$-algebras.
@@
We summarize the main contributions of the paper below. As the first contribution, we introduce the concept of partial law invariance as a new and relevant property for risk measures and preference relations in Section 2.1, with two substantive motivating examples in Sections 2.2 and 2.3, and a connection to stochastic dominance in Section 3.1. As the second contribution, a main theoretical result is presented, Theorem 1 in Section 3.2, which is a representation for coherent risk measures (Artzner et al. (1999)) satisfying partial law invariance. Third, generalizing the idea of a single source in Sections 23, the case of multiple sources is studied with a representation result Theorem 2 in Section 4, where we also explain the differences between the source theory of Baillon et al. (2023) and our approach. Fourth, we offer several extensions of the above results. A technical tool, called coherent adjustments, in Section 5.1 leads to an alternative representation of coherent risk measures in Theorem 3. This enables us to derive Theorem 4, which extends the well-known Kusuoka’s representation (Kusuoka (2001) and Frittelli and Rosazza Gianin (2005)) under a stronger notion of partial law invariance in Section 5.2. With no surprise to specialists in risk measures, most results can be smoothly generalized to the case of convex risk measures, and we present some of them in Section 5.3. Finally, formulas to compute the partially law-invariant ES are studied in Section 6, where we also present some applications and numerical examples. Section 7 concludes the paper. We will focus on an atomless probability space in the main part of the paper. The analysis in the case of a finite probability space is presented in the appendices, along with all proofs of technical results.
@@ 2.1 Risk measures, preference relations, and partial law invariance
We, throughout, fix a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\mathbb{P}$ is our reference probability model. We denote by $L^{0}$ the set of measurable functions, where functions that agree $\mathbb{P}$-almost surely (a.s.) are identified. A generic set $\mathcal{X}\subseteq L^{0}$ is interpreted as the collection of financial positions of interest, and $X\in\mathcal{X}$ represents the loss of the financial position for various states of the world $\omega\in\Omega$ [^1] A positive value of $X$ represents a loss, and negative value of $X$ represents a profit; this is the standard loss/profit convention in the risk measure literature. It will always be assumed that $\mathcal{X}$ contains the constant random variables, identified with elements of $\mathbb{R}$. A risk measure is a mapping $\rho:\mathcal{X}\to\mathbb{R}$, with the main interpretation of the amount of capital needed to make a financial loss $X$ acceptable.
@@
The property $\mathcal{F}$-law invariance is widely studied in the risk measure literature, called law invariance (or, probabilistic sophistication in decision theory). As far as we know, $\mathcal{G}$-law invariance with $\mathcal{G}\neq\mathcal{F}$ has not been studied in the literature. It is easy to see that law invariance is stronger than $\mathcal{G}$-law invariance. When making a contrast, we will refer to $\mathcal{G}$-law invariance as partial law invariance and to $\mathcal{F}$-law invariance as full law invariance, where the $\sigma$-algebras should be clear from the context. We also say that a set $\mathcal{X}\subseteq L^{0}$ is $\mathcal{G}$-law invariant if $X\in\mathcal{X}\cap L^{0}(\mathcal{G})$ and $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y\in L^{0}(\mathcal{G})$ imply $Y\in\mathcal{X}$.
@@
2.2 Robust risk assessment under model uncertainty
Suppose a firm’s portfolio manager is tasked with making investment decisions involving the financial positions in $\mathcal{X}$. Among the random variables in $\mathcal{X}$, there is some financial position $Z$ where the manager knows that the distribution of $Z$ is accurate under the probability model, i.e., there is no (or negligible) model uncertainty; for instance, $Z$ may be the losses from a regular property insurance portfolio, where there is enough data to confidently estimate its distribution. The distribution of the other random variables $X$ under the probability model may be subject to a considerable amount of model uncertainty, such as the losses of a stock without a sufficiently long history of data. The manager needs to decide on a risk measure. In this context, law invariance naturally applies when dealing with positions $g(Z)$ that are functions of $Z$ because the manager is confident about their distributions. However, for random variables $X$ that are not $\sigma(Z)$-measurable, if $X\buildrel\mathrm{d}\over{=}_{\mathbb{P}}g(Z)$, it is unreasonable to insist that $X$ and $g(Z)$ are equally risky due to the model uncertainty associated with $X$; the same can be said for two general random variables $X$ and $Y$. In other words, it is natural to require $\sigma(Z)$-law invariance but not (full) law invariance.
Denote by $\mathcal{M}_{1}$ the set of probability measures on $(\Omega,\mathcal{F})$ that are absolutely continuous with respect to $\mathbb{P}.$ Due to the presence of model uncertainty and to be robust to model misspecification, the manager may choose a collection of probability measures $\mathcal{Q}\subseteq\mathcal{M}_{1}$ with the property that for all $\mu\in\mathcal{Q}$, $F_{\mu,Z}=F_{\mathbb{P},Z}$, where $Z$ is the loss whose distribution the manager is certain about, and $F_{\mu,Z}$ is the cumulative distribution function (cdf) of $Z$ under $\mu$. For example, the manager could take
@@ 2.2 Robust risk assessment under model uncertainty
where $\epsilon>0$ and $d:\mathcal{M}_{1}\times\mathcal{M}_{1}\to[0,\infty]$is some measure of divergence, such as a Wasserstein distance or the Kullback–Leibler divergence. For $X\in\mathcal{X}$, the manager needs to evaluate the risk of $X$ under each $\mu\in\mathcal{Q}$, and then aggregate these evaluations in some way. For instance, if the manager uses the mean to quantify risk and the worst-case approach of Gilboa and Schmeidler (1989) to aggregate the risk evaluations, then the resulting risk measure is $\rho:X\mapsto\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X]$. We can see that $\rho$ is $\sigma(Z)$-law invariant because $\rho(X)=\mathbb{E}^{\mathbb{P}}[X]$ for $X$ that is $\sigma(Z)$-measurable.
Using the mean as a risk measure is overly simplistic. A risk measure common in financial practice is the Expected Shortfall (ES also known as CVaR). ES serves as the standard risk measure for market risk in the banking regulatory framework of BCBS (2019); see Rockafellar and Uryasev (2002) for its optimization and Wang and Zitikis (2021) for an axiomatization. Denote by $L^{\infty}$ the set of essentially bounded random variables on $(\Omega,\mathcal{F},\mathbb{P})$. For a given probability measure $\mu\in\mathcal{M}_{1}$, the ES at level $\alpha\in[0,1)$ is given by
$\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left\{x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}[(X-x)^{+}]\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
@@
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}_{\alpha}^{\mu}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$ (2)
@@
Another common way to assess model uncertainty is through the use of multiple prior probability measures. Law-invariant risk measures under multiple probability measures are studied by Wang and Ziegel (2021) under the name of scenario-based risk measures. It turns out that partial law invariance is closely related to scenario-based risk measures. Take a set $\mathcal{Q}\subseteq\mathcal{M}_{1}$. Wang and Ziegel (2021) defined the following property.
@@
This property means that if two risks are identically distributed under all probability measures of interest, then they are assessed as equally risky.
It is evident that law invariance is a particular case of both $\mathcal{G}$-law invariance, by choosing $\mathcal{G}=\mathcal{F}$, and $\mathcal{Q}$-basedness, by choosing $\mathcal{Q}=\{\mathbb{P}\}$. Moreover, law invariance is stronger in general than both $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness (if $\mathbb{P}\in\mathcal{Q}$). What is not so obvious is how $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness are connected, and we will first address this question. We denote by $L^{1}$ the set of integrable random variables on $(\Omega,\mathcal{F},\mathbb{P})$. For $\mu\in\mathcal{M}_{1}$, let $D_{\mu}=\mathrm{d}\mu/\mathrm{d}\mathbb{P}\in L^{1}$, that is, its Radon–Nikodym derivative with respect to $\mathbb{P}$.
Proposition 3.

### v5

# MNSC-0001-2024.00

@@
[…]
MNSC-0001-2024.00
[…]
Shen, Van Oosten, Wang
[…]
Partial Law Invariance and Risk Measures
[…]
Partial Law Invariance and Risk Measures
[…]
We introduce partial law invariance, generalizing law invariance and probabilistic sophistication widely used in decision theory, as well as statistical and financial applications. Partial law invariance may be interpreted as law invariance restricted to events for which there is no model uncertainty, reflecting practical situations in decision theory and financial risk management. We fully characterize partially law-invariant coherent risk measures via a novel representation formula. Strong partial law invariance is defined to bridge the gap between the above characterization and the classic representation formula of Kusuoka. We propose a few classes of new risk measures, including partially law-invariant versions of the Expected Shortfall and the entropic risk measures, and illustrate their applications in risk assessment under different types of uncertainty. We provide a tractable optimization formula for computing a class of partially law-invariant coherent risk measures and give a numerical example.
[…]
Law invariance, probabilistic sophistication, ambiguity, coherent risk measures, source theory, Expected Shortfall
@@ 1 Introduction
Law invariance is the fundamental property that makes risk measures, as well as many decision models, applicable in practice. Once a reference probability model is specified, either through statistical estimation or expert opinion, law-invariant risk measures evaluate financial positions that share the same distribution under this probability model equally. This connects the theory of risk measures to statistics, thus allowing for a description of the probabilistic structure of financial positions, arguably a basic step in any quantitative financial analysis. In decision theory, law invariance is often known as probabilistic sophistication (Machina and Schmeidler (1992)).
@@
Although law invariance is a fundamental concept, it may not be universally desirable in all contexts. For instance, model misspecification, or model uncertainty, occurs when the assumed probabilistic model for various financial positions does not accurately represent the true, unknown, distribution of these positions. This can be problematic as the value of a law-invariant risk measure may differ significantly between the assumed model and the objective truth. For instance, during the 2008 financial crisis, the probability model used for default rates of mortgages was questionable, and the loss distribution for mortgage-backed securities was underestimated, leading to insufficient capital requirements for many companies to remain solvent (e.g., McNeil et al. (2015)). Although we take a risk measure perspective in this paper, the problem of model uncertainty extends beyond this and can affect any (usually statistical) process that relies on estimating a probabilistic model, followed by computing a quantity based on that model. We will generally refer to uncertainty in the probabilistic model as “model uncertainty”, which is also known as Knightian uncertainty or ambiguity in the literature. There are many avenues of research that account for model uncertainty in risk assessment in different ways. One area is distributionally robust optimization, in which optimization with risk measures is performed against a worst-case scenario; see e.g., El Ghaoui et al. (2003), Natarajan et al. (2008), and Zhu and Fukushima (2009). Another area is to study risk measures formulated from multiple scenarios; see e.g., Delage and Li (2018), Delage et al. (2019) and Wang and Ziegel (2021). Model uncertainty is deeply rooted in decision theory, often called ambiguity in that literature. The goal there is to model preferences of decision makers who do not know the precise likelihood of events that govern their outcomes. Ambiguity is the key ingredient in the influential work of Choquet expected utility and maxmin expected utility pioneered by Schmeidler (1989) and Gilboa and Schmeidler (1989), followed by, e.g., the smooth ambiguity model of Klibanoff et al. (2005) and the variational preferences of Maccheroni et al. (2006).
Most of the literature, including the works cited above, addresses model uncertainty through three interrelated approaches. The first approach models financial positions as real-valued random variables (measurable functions), starting with a fixed reference probability model and a law-invariant risk measure, such as VaR, ES, or a function that represents expected utility preferences, that captures the risk preferences under full confidence in the reference model (that is, pure risk). Model uncertainty is then incorporated by aggregating the risk measure across a family of alternative probability models, often close to the reference probability in some way. This aggregation is typically the worst-case aggregation, as is common in distributionally robust optimization. The second approach, as in Artzner et al. (1999) and Föllmer and Schied (2016), also models financial positions as measurable functions, but it defines a risk measure directly on the space of these functions without fixing a reference probability model. This method handles model uncertainty by never supposing a reference probability model from the start, and it does not need a concept of pure risk (or preferences over pure risks). The third approach, common in decision theory, adopts the two-stage framework of Anscombe and Aumann (1963), by building preferences over state-dependent monetary lotteries. It begins with a law-invariant risk measure, typically one representing some expected utility preferences, but it is applied to the monetary lotteries, instead of the mappings from states to lotteries. Values of the risk measure over different states are aggregated in various ways such as in Schmeidler (1989), Gilboa and Schmeidler (1989), and Maccheroni et al. (2006). In this setting, pure risk refers to constant mappings that map each state to the same lottery.
@@
In this paper, we take a novel approach. We model financial positions as measurable functions and begin with a reference probability model but not a law-invariant risk measure. To account for model uncertainty, risk measures are assumed to satisfy law invariance only partially, that is, only with respect to random outcomes depending on events with no model uncertainty. This approach to modeling risk measures can address situations where some events are subject to more uncertainty than others. Traditional approaches, such as distributionally robust optimization, cannot handle this setting, as these models of uncertainty cannot directly incorporate sub-$\sigma$-algebras.
The following example can motivate this idea. After estimating a probability model for various loss factors, it is natural that different factors will have different levels of model uncertainty. For example, this difference might result from a statistical inference procedure, where the confidence regions for parameters tied to certain loss factors could be larger or smaller than those for other loss factors, for instance, due to data scarcity. From this point, we can typically divide the loss factors into two categories: the loss factors for which we have no, or negligible, model uncertainty and those that have non-negligible model uncertainty. We will call all the events related to the loss factors in the first category the “model-certain source.” A financial position is said to depend on this source if its outcome is determined by events within this source. The next step is to select a risk measure for decision making. Requiring law invariance would not be able to distinguish factors in or outside the model-certain source, which is not desirable. However, for financial positions depending on the model-certain source, it is reasonable to measure their risk level solely based on distributions given by the probability model, as these positions involve no model uncertainty; thus, a partial form of law invariance is desirable. With this guiding principle in mind, we introduce partial law invariance, based on a suitably chosen source, which is mathematically modeled by a sub-$\sigma$-algebra $\mathcal{G}$ of the whole model space. A risk measure is partially law invariant (with respect to $\mathcal{G}$) if it equates the risks of financial positions that (a) both belong to this source (i.e., $\mathcal{G}$-measurable) and (b) share identical distributions. This property reflects the discussion above with $\mathcal{G}$ being the model-certain source, but it is applicable beyond this interpretation.
@@
We summarize the main contributions of the paper below. As the first contribution, we introduce the concept of partial law invariance as a new and relevant property for risk measures and preference relations in Section 2.1, with two substantive motivating examples in Sections 2.2 and 2.3, and a connection to stochastic dominance in Section 3.1. As the second contribution, a main theoretical result is presented, Theorem 3.3 in Section 3.2, which is a representation for coherent risk measures (Artzner et al. (1999)) satisfying partial law invariance. Third, generalizing the idea of a single source in Sections 23, the case of multiple sources is studied with a representation result Theorem 4.1 in Section 4, where we also explain the differences between the source theory of Baillon et al. (2025) and our approach. Fourth, we offer several extensions of the above results. A technical tool, called coherent adjustments, in Section 5.1 leads to an alternative representation of coherent risk measures in Theorem 5.4. This enables us to derive Theorem 5.7, which extends the well-known Kusuoka’s representation (Kusuoka (2001) and Frittelli and Rosazza Gianin (2005)) under a stronger notion of partial law invariance in Section 5.2. With no surprise to specialists in risk measures, most results can be smoothly generalized to the case of convex risk measures, and we present some of them in Section 5.3. Finally, we discuss some explicit classes of partially law-invariant risk measures. Formulas to compute the partially law-invariant ES are studied in Section 6.1. To provide some concrete example, two constructions of partially law-invariant coherent risk measures are discussed in Section 6.2. An application with numerical examples is presented in Section 6.3. Section 7 concludes the paper. We will focus on an atomless probability space in the main part of the paper. The analysis in the case of a finite probability space is presented in the appendices, along with all proofs of technical results.
@@ 2.1 Risk measures, preference relations, and partial law invariance
We, throughout, fix a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\mathbb{P}$ is our reference probability model. We denote by $L^{0}$ the set of real-valued measurable functions, where functions that agree $\mathbb{P}$-almost surely (a.s.) are identified, and by $L^{\infty}$ and $L^{1}$ the subsets of random variables that are essentially bounded and integrable, respectively. A generic set $\mathcal{X}\subseteq L^{0}$ is interpreted as the collection of financial positions of interest, and $X\in\mathcal{X}$ represents the loss of the financial position for various states of the world $\omega\in\Omega$ [^1] A positive value of $X$ represents a loss, and negative value of $X$ represents a profit; this is the standard loss/profit convention in the risk measure literature. It will always be assumed that $\mathcal{X}$ contains the constant random variables, identified with elements of $\mathbb{R}$. A risk measure is a mapping $\rho:\mathcal{X}\to\mathbb{R}$, with the main interpretation of the amount of capital needed to make a financial loss $X$ acceptable.
@@
When $\mathcal{G}$ is the largest $\sigma$-algebra $\mathcal{F}$, the obtained $\mathcal{F}$-law invariance is the widely studied law invariance of the risk measure literature (or, probabilistic sophistication in decision theory). As far as we know, $\mathcal{G}$-law invariance with $\mathcal{G}\neq\mathcal{F}$ has not been studied in the literature. It is easy to see that law invariance is stronger than $\mathcal{G}$-law invariance. When making a contrast, we will refer to $\mathcal{G}$-law invariance as partial law invariance and to $\mathcal{F}$-law invariance as full law invariance, where the $\sigma$-algebras should be clear from the context. We also say that a set $\mathcal{X}\subseteq L^{0}$ is $\mathcal{G}$-law invariant if $X\in\mathcal{X}\cap L^{0}(\mathcal{G})$ and $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y\in L^{0}(\mathcal{G})$ imply $Y\in\mathcal{X}$.
@@
2.2 Partial law invariance and model uncertainty
In this section we elaborate on the role of partial law invariance in the context of model uncertainty.
Suppose a firm’s portfolio manager is tasked with making investment decisions involving the financial positions in $\mathcal{X}$. Among the random variables in $\mathcal{X}$, the manager knows that the distribution of a financial position $Z$ is accurate under the probability model $\mathbb{P}$; that is, there is no (or negligible) model uncertainty. For instance, $Z$ may be the losses from a regular property insurance portfolio, for which there are enough data to confidently estimate its distribution. The distribution of the other random variables $X$ under the probability model may be subject to a considerable amount of model uncertainty, such as the losses of a stock without a sufficiently long history of data. The manager needs to decide on a risk measure. In this context, law invariance naturally applies when dealing with positions $g(Z)$ that are functions of $Z$ because the manager is confident about their distributions. However, for random variables $X$ that are not $\sigma(Z)$-measurable, if $X\buildrel\mathrm{d}\over{=}_{\mathbb{P}}g(Z)$, it is unreasonable to insist that $X$ and $g(Z)$ are equally risky due to the model uncertainty associated with $X$; the same can be said for two general random variables $X$ and $Y$. In other words, it is natural to require $\sigma(Z)$-law invariance but not (full) law invariance.
Denote by $\mathcal{M}_{1}$ the set of probability measures on $(\Omega,\mathcal{F})$ that are absolutely continuous with respect to $\mathbb{P}.$ The manager may choose a collection of probability measures $\mathcal{Q}\subseteq\mathcal{M}_{1}$ to reflect model uncertainty. Because $Z$ is the loss whose distribution the manager is certain about, the set $\mathcal{Q}$ should satisfy that for all $\mu\in\mathcal{Q}$, $F_{\mu,Z}=F_{\mathbb{P},Z}$, where $F_{\mu,Z}$ is the cumulative distribution function (cdf) of $Z$ under $\mu$. For example, the manager could take
@@ 2.2 Partial law invariance and model uncertainty
where $\epsilon>0$ and $d:\mathcal{M}_{1}\times\mathcal{M}_{1}\to[0,\infty]$ is some measure of divergence, such as a Wasserstein distance or the Kullback–Leibler divergence. For $X\in\mathcal{X}$, the manager needs to evaluate the risk of $X$ under each $\mu\in\mathcal{Q}$, and then aggregate these evaluations in some way. For instance, if the manager uses the mean to quantify risk and the worst-case approach of Wald (1945, 1949) (axiomatized by Gilboa and Schmeidler (1989)) to aggregate the risk evaluations, then the resulting risk measure is $\rho:X\mapsto\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X]$. We can see that $\rho$ is $\sigma(Z)$-law invariant because $\rho(X)=\mathbb{E}^{\mathbb{P}}[X]$ for $X$ that is $\sigma(Z)$-measurable.
Using the mean as a risk measure is overly simplistic. A risk measure common in financial practice is the Expected Shortfall (ES also known as CVaR). ES serves as the standard risk measure for market risk in the banking regulatory framework of BCBS (2019); see Rockafellar and Uryasev (2002) for its optimization and Wang and Zitikis (2021) for an axiomatization. For a given probability measure $\mu\in\mathcal{M}_{1}$, the ES at level $\alpha\in[0,1)$ is given by
$\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}[(X-x)^{+}]\right),~{}~{}~{}~{}X\in L^{\infty}.$
@@
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}_{\alpha}^{\mu}(X),~{}~{}~{}~{}X\in L^{\infty}.$ (2)
@@
Another common way to assess model uncertainty is through the use of multiple prior probability measures. Law-invariant risk measures under multiple probability measures are studied by Wang and Ziegel (2021) under the name of scenario-based risk measures, which are formulated with the following property, for a set $\mathcal{Q}\subseteq\mathcal{M}_{1}$.
@@
This property means that if two risks are identically distributed under all probability measures of interest described by $\mathcal{Q}$, then they are assessed as equally risky.
It is evident that law invariance is a particular case of both $\mathcal{G}$-law invariance, by choosing $\mathcal{G}=\mathcal{F}$, and $\mathcal{Q}$-basedness, by choosing $\mathcal{Q}=\{\mathbb{P}\}$. Moreover, law invariance is stronger in general than both $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness (if $\mathbb{P}\in\mathcal{Q}$). What is not so obvious is how $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness are connected, which we address next. For $\mu\in\mathcal{M}_{1}$, let $D_{\mu}=\mathrm{d}\mu/\mathrm{d}\mathbb{P}\in L^{1}$, that is, its Radon–Nikodym derivative with respect to $\mathbb{P}$.
Proposition 2.3